OpenStudy (anonymous):

cos θ - sin θ = √2 sin θ/2

OpenStudy (anonymous):

must be between the interval \[[0,2\pi]\]

OpenStudy (anonymous):

the directions say this, "Use a Double- or Half-Angle Formula to solve the equation in the interval [0,2pi]." The answers in the book say \[\pi/6,3\pi/2\]

OpenStudy (jdoe0001):

\(\bf cos(\theta)-sin(\theta)=\cfrac{\sqrt{2}\quad sin(\theta)}{2}\quad ?\)

OpenStudy (anonymous):

thats what it appears to be naturally just the line isn't there but yes thats the problem :)

OpenStudy (anonymous):

I remember my professor telling me that in problems like this, to always do the side that looks more ocmplicated

OpenStudy (anonymous):

One sec let me google search the half angle and double angle formulas

OpenStudy (anonymous):

I tried dividing out the \[\sqrt{2}\] but that didnt help, and converting the sin in to a square root but it didnt help. Gotta love College Trig classes I like calculus better :)

OpenStudy (anonymous):

the book says the answers are \[\pi/6,3\pi/2\]

OpenStudy (anonymous):

I'm working it out right now

OpenStudy (anonymous):

Sorry I'm taking so long I just seem to be having a hard time finding the half and souble angle formulas

OpenStudy (anonymous):

i got them handy :) ill type them in

OpenStudy (anonymous):

And are you sure it's not....

OpenStudy (anonymous):